OpenStudy (anonymous):

The part of the lecture talking about GPS, shouldn't it be S/h to get L? Why is it L/h?

OpenStudy (phi):

Can you post a link to the lecture ?

OpenStudy (phi):

In the GPS example, the professor did not use S at all, so S/h is not defined. The point of his discussion (which may be lost amongst the discussion of the GPS), is this: Let's say you have a variable that is the function of another variable. Most often, we might say y = f(x) though in this case, we say L= f(h) Now the question we are interested in, is if we put in \(x +\Delta x\) rather than x, how much will L change? (The professor associated Δx with an unknown error) In other words, what is the difference \(\Delta L\)= f( \(x +\Delta x\)) - f(x) He went on to say \[ \frac{\Delta L}{\Delta x} = \frac{f(x+\Delta x)-f(x)}{\Delta x} \approx \frac{dL}{dx}\] (only in the limit as Δx-> 0, can we equate the expressions) He then said that the derivative is easy to find, if we have the function. For example, if the satellite is at a known altitude A, we can use the pythagorean theorem to solve for L in terms of the fixed A and the measured h: \[ L = \sqrt{h^2 -A^2} \] we use calculus to find dL/dh to be (I don't think you learned this part yet) \[\frac{dL}{dx}= \frac{h}{\sqrt{h^2 -A^2}}\] The main point is that this expression tells us the ratio of the change in L caused by a change in x. If the change in x is "small", we can say \[ \frac{\Delta L}{\Delta x} \approx \frac{dL}{dx} \] and, more usefully \[ \Delta L \approx \frac{dL}{dx} \Delta x \] or, if we use the actual derivative \[ \Delta L \approx \frac{h}{\sqrt{h^2 -A^2}} \Delta x \] that is a useful expression. (though perhaps you don't yet see why)